Question: Steve traveled 600 miles at a certain speed. Had he gone 20mph faster, the trip would have taken 1 hour less. Find the speed of his vehicle.
Here's what i've done, but i'm not sure if it is correct.
d=rt
600=(x-1)20
600=20x-20
+20 +20
620=20x (i divided both sides by 20)
31=x

Answer by scott8148(6628)

"Steve traveled 600 miles at a certain speed" __ 600=rt __ 600/r=t

"Had he gone 20mph faster, the trip would have taken 1 hour less" __ 600=(r+20)(t-1) __ 600=rt+20t-r-20

substituting __ 600=(600)+20(600/r)-r-20 __ subtracting 600 __ 0=12000/r-r-20

multiplying by -r __ 0=r^2+20r-12000 __ factoring __ 0=(r+120)(r-100)

0=r+120 __ -120=r __ negative value not realistic

0=r-100 __ 100=r __ heavy foot